OpenStudy (anonymous):

solve each equation check your answer 7j=28

OpenStudy (chrissy):

divide both sides by 7

OpenStudy (chrissy):

j=4

OpenStudy (anonymous):

j= 28/7 j=4

OpenStudy (anonymous):

where did you get the 7 to divided by

OpenStudy (anonymous):

ok 7j = 28 7/7 j = 28/7 whats 7/7= 1 and 28/7 = 4 so j* 1 = 4 or you can say j = 4